A circular current–carrying coil has a radius R. Determine the distance in terms of R, of a point on the axis from the centre of the coil where the magnetic field produced by the coil is 1/8 of its value at the centre of the coil.


1
Expert's answer
2022-01-23T16:16:19-0500

"B(z)=\frac{\mu_0}2\frac{IR^2}{(z^2+R^2)^{\frac 32}},"

"\frac{B_0}{B_1}=\frac{(z^2+R^2)^{\frac 32}}{R^3}=8,"

"z=\sqrt3R."